Mir Extension Update 12-16 June 1995

MIR EXTENSION OPERATIONS WEEKLY REPORT FOR WEEK OF Monday, 12.06.95 through Friday, 16.06.95
Report filed by Sarah Kirby

MIR OPERATIONS

Monday (6/12):

The crew spent the day at rest. They each enjoyed a private conference with their families. A few medical experiments and some minor EVA preparations were also conducted.

The crew burned three O2 candles today. Late in the day, the commands to activate the Kvant-2 array articulation drive were read up. The crew was told that the last time the drive was activated a unit started to overheat. The crew was asked to monitor that unit and if it looked like it was overheating to tell the ground because they had some plans to provide fans for more active cooling of the unit.

Tuesday (6/13):

The comm session through Dryden did not work. Dryden reported a good uplink, but no downlink (not even static) was received so it is entirely possible that the crew chose not to participate. Later in the day Soloviev cancelled all American ground pass attempts until at least next Monday.

There is a problem in the packet transmission from both the ground and the crew. Therefore the crew reported on the open loop a problem that occurred during the last move of Kristall (the core module shook). The crew also reported that one of the spare suits (which they are evidently preparing for the EVA) "stinks". They had already been blowing air through it for two days.

Soloviev asked the crew to do their best to remove the TREK detector on the next EVA if time permits at the end. Later the crew said that the pressure was "up to the mark" and that they were feeling much better. Finally, the crew did a test on the Strela to see if they could tell whether it was hindering the Kvant-2 array. The crew said they thought that the boom might be hindering the array but that the EVA should provide a definite answer.

Wednesday(6/14):

The crew had a two-way video conference with the ground. The primary focus of the discussion was the Spektr array. The crew showed the ground a sketch of what they see. The ground showed the crew their sketch as well as some video of a dry hydrolab test that was done to try and develop a procedure that might work. The crew also showed more taped video of the Spektr array, the folded Kristall array, and some EVA tool ideas for the Spektr EVA.

There was quite a bit of discussion about the length of the Strela boom. The crew seems to think that it is fully extended and that it is two meters shorter than what the ground thinks. The ground thinks that perhaps two segments on the boom are not fully extended, but the crew is not accepting this explanation.

The packet transmission problem was finally solved when the crew turned on an amplifier.

A Russian science test called Astra is scheduled for the 19th-22nd which will involve the use of the TORU equipment to fire some thrusters to do ACS plume environment analyses. I presume that they ve already modelled the rest of the station and that only the Spektr plumes have yet to be characterized. It seems like the TORU is the easiest way to command only Spektr jets to fire. After the test the crew will disassemble the TORU equipment for return on STS-71. The crew made some comment about being concerned that the test could adversly impact some cables due to plume impingement.

Thursday(6/15):

The crew began the day talking EVA preparation medical tests. Contrary to the plan as we left it the night before, no video was exchanged. Several Q&A periods were conducted over the course of the day clarifying various aspects of the EVA plan.

During one conversation, Soloviev said that one time in the past they had a similar length problem with the Strela. When the crewman tried one last time, he was able to extend the boom another 1/2 meter. The crew still insisted that the boom was fully extended. There was some talk about checking some "cones" on the Kristall and Spektr. The timeline made it look like these were some sort of instrumentation cones.

The crew also reported that after the Spektr array deploy, they saw the protective cover that had kept the arrays safe during ascent follow the Mir for a couple of revs.

They talked about doing the Astra experiment without using the Spektr thrusters (so evidently the purpose is not solely to characterize plume environments). The main concern is plume impingement on some cables. The only thing that I can think of is that there might be some external cables on Kristall (part of the power transfer cables after the array move?). There shouldn t be any cables of concern on Spektr itself since the thrusters were used during Spektr s free flight and rendezvous. The only thrusters in our drawing are located at the base of Spektr so Kristall interference is plausible.

In the few minutes before the CNBC press conference, the crew was told that the chief designer of the arrays was not satisfied with the situation and so the EVA was cancelled. Tools will be manufactured and possibly flown on STS-71 so that the Mir-19 crew can do the EVA to fix the array. The crew was told that launch of STS-71 was scheduled for 6/23 or 6/24.

During the press conference Norm said that the crew barely felt Spektr (or the Progress) dock and that they expected a similar sensation for the Shuttle dock.

Friday(6/16):

The crew had a light day. Mostly they rested. They also spent some time putting away EVA equipment and the EVA suits. The ground told the crew that they had discovered a way to get another 13 Amps from Spektr to the base block, but they didn t go into any details.

A radiogram was uplinked to the crew with some of the return-to-earth items to be packed for STS-71. One of the items the crew discussed was the Salyut-5B computer. The Salyut-5B is in Kristall, behind panel 205. This means that its not an MCS computer, but one of the computers in the OCCS. They also talked about a circuitboard for channel A of another Salyut computer that had previously been removed that they were to find for return as well as a cable for the computer.

MIR BALLISTICS: Beta angle changed from 71.3 to 74.9 degrees

DAY # of Altair Scheduled Sessions
Monday (6/12) 1
Tuesday (6/13) 0
Wednesday (6/14) 1 (2 way TV)
Thursday (6/15) 3
Friday (6/16) 7 (6 for EVA were cancelled)

MIR MAINTENANCE PLANNING

On Tuesday I met with Mr. Lapin who is Mr. Prokopiev s designated deputy while he is on vacation. We managed to discuss all of the items "above the line" from my fax. Our next meeting will be during STS-71 and will focus on plans for our discussions after the MEAT leaves Moscow.

On Wednesday I talked to Yuri Skurski and the motion control engineer. Ryumin has not yet commented on the letter to his subordinates because he was on vacation until earlier this week. Therefore, Yuri was hesitant to give me what is in fact design and not maintenance information until he gets a reading on Ryumin s interpretation of the letter. I chose not to press the issue at this time. Yuri and the MCS answered all of the questions that I had on that system since they were related to maintenance operations and responses to off-nominal Mir systems operations.

MEAT OPERATIONS COMMENTS:

Tuesday (6/13):

The flight surgeon was able to ask the life support person (Elena) some questions about the cabin pressure variation after she responded to the crew s query that it wasn t a problem. Elena is attributing the extreme variations that we see on the ground primarily to ground site variations (they have different cal curves?), especially Petropovlask (PPK) which usually has the lowest readings. The on-board manometer is accurate to within +/- 1 mmHg but is not in the telemetry. The crew readings on this have only varied +/- 5 mmHg or so. Then Elena pulled out a plot beginning on flight day 78 on top of which she overlaid the O2 cartridge burning and Elektron activation data. The plot evidently showed a steady variation with no indications of a leak before/after the Kristall move to/from the -Z.

The flight surgeon also reported that the hydrolab in Star City is dry, but they tried a Spektr access procedure anyway -- with great difficulty. It changed several people s minds about the difficulty of what they were asking the crew to do.

On the pressure situation, 667mmHg is the maximum cabin pressure at docking. The pressure will cycle around this by about +/-10 mmHg. After the EVA the maximum pressure will probably be slightly less. The only way to increase the pressure is to get a Progress resupply flight. There is an emergency reserve of air in some tanks, but it can not be refilled so they don t want to use it for a non-emergency purpose. They were designed to repress a Transfer compartment with an open hatch so that the crew could ingress.

  • submit to reddit


Loading

Kepler Communications - Aether - Unbound Connectivity to your on-orbit assets
Stolen Skies by Tim Powers - Baen Books
Don’t Blow Yourself Up - The Further True Adventures and Travails of the Rocket Boy of October Sky By Homer Hickam




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on December 16, 1995 4:52 PM.

Russian Doubts About The Future in Space was the previous entry in this blog.

NASA Staff Meeting Minutes 16 January 1996 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.